OpenStudy (anonymous):

If 2x = 5, 3y = 4, and 4z = 3, what is the value of 24xyz ? a.12 b.18 c.36 d.48 e.60

OpenStudy (anonymous):

e.60

OpenStudy (anonymous):

\[2x \times 3y \times 4z = 5 \times 4 \times 3 \rightarrow 24xyz=60\] Hence option e is correct

OpenStudy (anonymous):

I still don't get why you would assume the equation of 2x x 3y x 4z is pertinent to anything. and what does the 24 have to do with anything. In a standard algebra wouldn't you substitute x=5/2 and y=4/3 and z=3/4 and multiply that all by the 24 to get a number just a little shy of 10?

OpenStudy (anonymous):

Since you had to find the value of 24xyz and not of xyz. That's why i multiplies 2x x 3y x 4z and did not substitute x=5/2 and y=4/3 and z=3/4 in the problem.

OpenStudy (anonymous):

Sorry, I'm not getting any logic behind this one.